OpenStudy (anonymous):

what number mulitplied by the numerator and added to the denominator of 4/9 makes the resulting fraction equilavent to 7/4?? the number is?

OpenStudy (anonymous):

@ArkGoLucky

OpenStudy (anonymous):

um I don't think this is possible because the numerator of the first fraction is not changed so it can't equal the second fraction. ever

OpenStudy (anonymous):

4 does not equal 7

OpenStudy (anonymous):

it dosent' have a answer box for not a answer

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

I have no idea either

OpenStudy (anonymous):

oh wait nvm, there is an answer. The number just has to be a fraction

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

Is there a box for a fraction?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

okay cool

OpenStudy (anonymous):

do u know it? lol

OpenStudy (anonymous):

okay got

OpenStudy (anonymous):

it

OpenStudy (anonymous):

so here's what you do so make an equation say x is the number you're looking for try to think of a way you can make an equation using x

OpenStudy (anonymous):

ask if you need help

OpenStudy (anonymous):

I'll give you a hint use 4x

OpenStudy (anonymous):

ok so what do i do now with the 4x?

OpenStudy (anonymous):

is the answer going to be 2? @ArkGoLucky

OpenStudy (anonymous):

not quite. I'll just show you the equation. So you know that the number x, when multiplied by the numerator(4) then added to the denominator(9) equals 7/4 That means 4/9+4x=7/4

OpenStudy (anonymous):

so 7/4 is the answer or do i solve that?

OpenStudy (anonymous):

no, now you solve for x, which is the number you're looking for

OpenStudy (anonymous):

i got 47/144?

OpenStudy (anonymous):

is that right or wrong?

OpenStudy (anonymous):

almost

OpenStudy (anonymous):

the numerator is right except it's a negative and the denominator is wrong

OpenStudy (anonymous):

how did i get the denominator wrong?

OpenStudy (anonymous):

did you solve x in\[\frac{ 4 }{ 9+4x }=\frac{ 7 }{ 4 }\]

OpenStudy (anonymous):

yea and i got that big fraction

OpenStudy (anonymous):

try plugging in your answer into that equation

OpenStudy (anonymous):

okay I'll show you how to solve for x multiply both sides of the equation by the common denominator. In this case, it is (4)(9+4x) this is to rid of the fractions

OpenStudy (anonymous):

when you do this, you get 16=7(9+4x) I'm sure you can solve for x now